Ligand A and protein B bind to each other with a Kd of 15 μM. If the total concentrations of A and B are 456 μM and 1,060 μM, respectively, then what fraction of A is bound?
Answer #1

In a bimolecular equilibrium reaction  

[A] + [B] ==> [AB], Ka = [AB] / [A][B]

[AB] ==> [A] +[B]

Kd = [A][B]/[AB]

From this the fraction of A Bound = [AB] / [A] + [AB] = 1/ ([A]/[AB] + 1) = 1/(Kd/[B] +1) = [B] / Kd + [B]

[A] (ligand) = 456 uM

[B] (protein) = 1060 uM

Kd = 15 uM

Fraction of A Bound = 1060 uM /(15 + 1060) uM = 0.9860

Fraction of Ligand A bound to protein = 0.9860.